fuyo:

The sides of a square are 3 2/7 inches long. What is the area of the square?

supie:

Do you mean 3^2/7 or no?

ellapayne:

wait are you supposedto explain out how you have the answer?

ellapayne:

supposed to*

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ellapayne wait are you supposedto explain out how you have the answer? \(\color{#0cbb34}{\text{End of Quote}}\) yes

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Do you mean 3^2/7 or no? \(\color{#0cbb34}{\text{End of Quote}}\)

fuyo:

yeas

fuyo:

1 attachment
supie:

Ok so, side=3^2/^7 We don't know the area yet obviously..... so, We have to use the formula: Area= side x side Area= 3^2/^7 Area=? i think

fuyo:

brb again

supie:

\(\color{#0cb}{\text{Originally Posted by}}\) @supie Ok so, side=3^2/^7 We don't know the area yet obviously..... so, We have to use the formula: Area= side x side Area= 3^2/^7 Area=? i think \(\color{#0cb}{\text{End of Quote}}\)

fuyo:

ooooh

fuyo:

1 attachment
fuyo:

das the options